A charge q1= 7.00 µC is located at the origin, and a second charge q2 = -5.00 µC is located on the x axis, 0.30 m from the origin. (A) Find the electric field at the point P, which has coordinates (0, 0.400) m. (B) If a charge q3 = -3.00 µC is placed at point P, what is the coulomb force experienced by this charge due to other charges q1 and q2. 


1
Expert's answer
2021-06-22T17:00:09-0400

Gives

"q_1=7.00\mu C;q_2=-5.00\mu C"


"r_1=0.40m;r_2=\sqrt{.40^2+.30^2}=0.50m"

"sin\theta=\frac{0.40}{0.50}=\frac{4}{5}"

"cos\theta=\frac{0.30}{0.40}=\frac{3}{4}"


"E_1=\frac{kq_1}{r_1^2}=\frac{{9\times10^9}\times7\times10^{-6}}{0.40^2}=3.9310^{5}N/C"

"E_2=\frac{kq_2}{r_2^2}=-\frac{{9\times10^9}\times5\times10^{-6}}{0.50^2}=1.8\times10^{5}N/C"

"E_2" x-axis components


"E_2cos\theta"

"E_2" y-axis components

"-E_2sin\theta"

"E_1=3.93\times10^5\hat{j}N/C"


"E_2=1.08\hat{i}+{(-1.44)}\hat{j}\times10^{5}N/C"

"E=E_1+E_2"


"E=1.1\times10^5\hat{i}+2.5\times10^{5}\hat{j} N/C"

Force

"q_3=-3.0\mu C"

"F_{13}=\frac{Kq_1q_3}{r_2^2}"


"F_{13}=-\frac{{9\times10^9}\times7\times10^{-6}\times3\times10^{-6}}{0.50^2}=0.756N"

"F_{23}=\frac{Kq_2q_3}{r_1^2}"


"F_{23}=\frac{{9\times10^9}\times5\times10^{-6}\times3\times10^{-6}}{0.40^2}=0.843N"